### v1

# Boltzmann priors for Implicit Transfer Operators

Abstract
Accurate prediction of thermodynamic properties is essential in drug discovery and materials science. Molecular dynamics (MD) simulations provide a principled approach to this task, yet they typically rely on prohibitively long sequential simulations. Implicit Transfer Operator (ITO) Learning offers a promising approach to address this limitation by enabling stable simulation with time steps orders of magnitude larger than MD. However, to train ITOs, we need extensive, unbiased MD data, limiting the scope of this framework. Here, we introduce Boltzmann Priors for ITO (BoPITO) to enhance ITO learning in two ways. First, BoPITO enables more efficient data generation, and second, it embeds inductive biases for long-term dynamical behavior, simultaneously improving sample efficiency by one order of magnitude and guaranteeing asymptotically unbiased equilibrium statistics. Further, we showcase the use of BoPITO in a new tunable sampling protocol interpolating ITO models trained on off-equilibrium simulation data and an unbiased equilibrium distribution to solve inverse problems in molecular science.
1 Introduction
Efficient molecular dynamics (MD) simulation on long time-scales is critical to a large number of scientific and engineering applications. Stable simulations rely on solvers taking tiny integration time-steps, making the simulation of most phenomena impractical with current methods. Since these simulations are stochastic, a simulation step corresponds to drawing a sample from a transition probability density $p(\mathbf{x}_{t+\tau}\mid\mathbf{x}_{t})$, where $\tau$ is a tiny time-step. Recently, deep generative models have emerged as a promising strategy to potentially speed up these simulations by learning transition probability densities where $\tau$ is much larger (Schreiner et al., 2023; Klein et al., 2023; Hsu et al., 2024; Fu et al., 2023) and thereby allow efficient sampling at long time-scales. Implicit Transfer Operator (ITO) learning (Schreiner et al., 2023) learns such surrogate models at multiple time-resolutions. While ITO has shown promise in accelerating simulations, it relies on extensive unbiased simulation data which may not always be available. In Markov state models, this limitation can be mitigated by integrating off-equilibrium simulations and enhanced sampling simulations (Trendelkamp-Schroer et al., 2015; Rosta & Hummer, 2014; Wu et al., 2014; 2016). However, such estimators are so far unavailable for deep generative surrogates of the transition density.
Figure 1: Boltzmann Priors for Implicit Transfer Operators (BoPITO) leverage pre-trained Boltzmann Generators to enable data-efficient training of surrogate models of the transition density after $N$ simulation steps. BoPITO is implemented using Score-Based Diffusion Models. The score of the model, $s_{\mathbf{\theta}}$, separates the contributions from the first eigenfunction of the transfer operator (equilibrium density), $s_{\text{eq}}$, from the rest, $s_{\text{dyn}}$. BoPITO embeds inductive biases for long-term dynamical behavior and enables interpolation between off-equilibrium and equilibrium models.
Here, we introduce Boltzmann Priors for Implicit Transfer Operator (BoPITO) learning (Figure 1). BoPITO leverages pre-trained Boltzmann Generators (BG) (Noé et al., 2019; Viguera Diez et al., 2024; Klein & Noé, 2024; Köhler et al., 2023; 2020; Midgley et al., 2024; 2023) to enable data-efficient training of ITO models, leading to one order of magnitude reduction in the simulation data needed for training. As the BG encodes the invariant measure or Boltzmann distribution of the dynamics encoded by the transition density, BoPITO, by construction, guarantees asymptotically unbiased equilibrium statistics. In this way, BoPITO can combine off-equilibrium data and biased data encoded into a BG prior to train ITO models that predict MD across multiple time-scales. Using BoPITO we introduce a new sampling strategy to recover approximate dynamics from biased off-equilibrium data, a BG prior, and unbiased time-correlation data, providing a new method for inverse problems for molecular systems.
Our main contributions are
  1. 1.
    Boltzmann Priors for Implicit Transfer Operator Learning (BoPITO): We provide a principled way to leverage Boltzmann Generators as priors for Implicit Transfer Operator Learning, boosting sample efficiency, ensuring the generation of uncorrelated data and allowing asymptotically unbiased equilibrium statistics for $\tau\rightarrow\infty$.
  2. 2.
    BoPITO interpolators: A tunable sampling protocol facilitating interpolation of BoPITO models trained on off-equilibrium simulation data to an unbiased equilibrium distribution.
  3. 3.
    We show that BoPITO interpolators can recover approximate dynamics from models trained on biased simulations. We frame the optimization of interpolation parameters as an inverse problem and select the interpolated ensemble that is most consistent with unbiased observables.
2 Background and preliminaries
2.1 Molecular dynamics and observables
Molecular dynamics (MD) is a widely used simulation method in chemistry, physics, and biology. It combines a mathematical model of the dynamics — e.g. Langevin dynamics (Langevin, 1908) — with a potential energy model, $U(\mathbf{x}):\Omega\rightarrow\mathbb{R}$, of a system of interest, in turn providing detailed mechanistic insights of molecular systems, through the time evolution of particles in configuration space, $\mathbf{x}\in\Omega$. Practically, MD is solved by numerical integration, and time is discretized, with step $\tau$. Consequently, we can understand MD as a Markov process with a Normal transition density $p(\mathbf{x}_{t+\tau}\mid\mathbf{x}_{t})$, whose associated Markov operator has the Boltzmann distribution,
$\mu(\mathbf{x})=\mathcal{Z}^{-1}\exp(-\beta U(\mathbf{x})),\text{with }\mathcal{Z}=\int\mathrm{d}\mathbf{x}\,\exp(-\beta U(\mathbf{x})),$ (1)
as its invariant measure or stationary distribution, where $\beta$ is the inverse temperature.
One important application of MD is to compute expectations or observables (Olsson, 2022):
  1. 1.
    Stationary observables:
    $O_{a}=\mathbb{E}_{\mu}\left[a(\mathbf{x})\right].$ (2)
  2. 2.
    Dynamic observables / Time correlation functions:
    $O_{a(t),b(t+N\tau)}=\mathbb{E}_{\mathbf{x}_{t}\sim\mu}\left[\mathbb{E}_{\mathbf{x}_{t+N\tau}\sim p_{\tau}(\mathbf{x}_{t+N\tau}|\mathbf{x}_{t})}\left[a(\mathbf{x}_{t})\cdot b(\mathbf{x}_{t+N\tau})\right]\right],$ (3)
where $p_{\tau}(\mathbf{x}_{t+N\tau}|\mathbf{x}_{t})$ is the conditional probability density after $N$ simulation steps with time-step $\tau$. The maps $a$ and $b:\Omega\rightarrow\mathbb{R}^{L}$ serve as observable functions or ‘forward models’ characterizing microscopic observation processes, e.g. indicating whether a drug is bound or not, or an interatomic distance, leading to observables such binding affinities and off-rates of a drug to a target protein.
Unfortunately, stable MD simulations rely on using time discretizations, $\tau$, on the order of $10^{-15}\,\mathrm{s}$, whereas properties such as protein folding or ligand unbinding occur on much longer time scales, $\sim 10^{-3}-10^{-1}\,\mathrm{s}$. Due to the sequential nature of these simulations, an impractical number of simulation steps are necessary to predict these properties in an unbiased fashion. As a result, most MD data will be ‘off-equilibrium’, e.g. trajectories exploring one or a few of the modes of $\mu(\mathbf{x})$.
2.2 Boltzmann Generators
Boltzmann Generators (BG) (Noé et al., 2019) are a generative machine learning approach to draw samples i.i.d. from the Boltzmann distribution, $\mu(\mathbf{x})$, of physical many-body systems (eq. 1). BGs learn a diffeomorphic map, $\mathcal{F}_{\bm{\theta}}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$, from a latent space, equipped with a simple base density $p(\mathbf{z})$, to configurational space of a physical system, such that the push-forward density $\hat{p}(\mathbf{x})=\mathcal{F}_{\bm{\theta}}\#p(\mathbf{z})$ closely approximates the Boltzmann distribution (eq. 1). In practice, $\mathcal{F}_{\bm{\theta}}$ is implemented using an invertible neural network architecture with tractable Jacobian determinants to enable efficient sampling and exact sample likelihood computation (Chen et al., 2018; Papamakarios et al., 2021). BGs are trained either with approximately equilibrated simulation data or with biased simulations, from i.e. enhanced sampling, by employing appropriate reweighting (Ferrenberg & Swendsen, 1989; Shirts & Chodera, 2008). Unbiased samples can then be generated by importance re-sampling. Similarly, unbiased expectations can be computed using the importance weights $w(\mathbf{x})=e^{-\beta U(\mathbf{x})}/\hat{p}(\mathbf{x})$. As such, BGs are surrogates of the target Boltzmann distribution, but do not model time-correlation statistics.
2.3 Transfer operators
Transfer operators (Ruelle, 1978; Schütte et al., 2009) provide a framework to describe the evolution of probability densities over time. Let $p$ denote an initial probability density function on $\Omega$, and $\rho$ its $\mu$-weighted version, $p=\mu\rho$. The Markov operator, $T_{\Omega}$, is defined using a transition density $p(\mathbf{x}_{t+\tau}|\mathbf{x}_{t})$,
$[T_{\Omega}(\tau)\circ\rho](\mathbf{x}_{t+\tau})\equiv\frac{1}{\mu(\mathbf{x}_{t+\tau})}\int_{\Omega}\mu(\mathbf{x}_{t})\rho(\mathbf{x}_{t})p(\mathbf{x}_{t+\tau}|\mathbf{x}_{t})\,\mathrm{d}\mathbf{x}_{t}.$ (4)
This operator describes the $\mu$-weighted evolution of absolutely convergent probability density functions on $\Omega$ by a discrete-time increment $\tau$ given by the dynamics encoded in a transition density $p(\mathbf{x}_{t+\tau}|\mathbf{x}_{t})$. In the context of molecular dynamics, $T_{\Omega}$ is a $\mu$-weighted equivalent to the Markov operator discussed above (sec. 2.1). The spectral form of the transfer operator is expressed as
$[T_{\Omega}(\tau)\circ\rho](\mathbf{x}_{t+\tau})=\sum_{i=1}^{\infty}\lambda_{i}(\tau)\langle\rho|\phi_{i}\rangle\ \psi_{i}(\mathbf{x}_{t+\tau}),$ (5)
where $\lambda_{i}(\tau)=\exp(-\tau\kappa_{i})$ are the eigenvalues, $\langle f|g\rangle=\int_{\Omega}f(\mathbf{x})g(\mathbf{x})\,\mathrm{d}\mathbf{x}$ and $\psi_{i}$ and $\phi_{i}$ are the corresponding right and left eigenfunctions, respectively. The eigenvalues $\lambda_{i}(\tau)$ depend on $\tau$, and $\kappa_{i}$ are characteristic relaxation rates. When the dynamics are reversible, all eigenvalues $\lambda_{i}$ are real-valued and lie in the interval $-1<\lambda_{i}\leq 1$. Notably, there is one eigenvalue $\lambda_{1}=1$ and its associated eigenfunction corresponds to the stationary distribution $\mu(\mathbf{x})$. Moreover, the left and right eigenfunction are related by a factor of the stationary density, $\phi_{i}(\mathbf{x})=\mu(\mathbf{x})\psi_{i}(\mathbf{x})$, and $\psi_{1}(\mathbf{x})=\mathbf{1}$.
2.4 Diffusion models
Diffusion models (DMs) (Ho et al., 2020; Song et al., 2020) are a popular generative modeling framework that approximates data densities $p(\mathbf{x}^{0})$ by learning to invert a noising process (‘forward diffusion’). The forward diffusion process is pre-specified and incrementally transforms the data distribution into a simple prior distribution, $p(\mathbf{x}^{T})$, through simulation of a time-inhomogenous Markov process represented by the SDE,
$\mathrm{d}\mathbf{x}^{t}=f(\mathbf{x}^{t},t)\,\mathrm{d}t+g(t)\,\mathrm{d}W,$ (6)
where ($0<t<T$) is the diffusion time, $f$ and $g$ are chosen functions, and $\mathrm{d}W$ is a Wiener process. To generate samples from the data distribution, $p(x^{0})$, we can sample from $p(\mathbf{x}^{T})$ and solve the backward diffusion process (‘denoising process’) (Anderson, 1982),
$\mathrm{d}\mathbf{x}^{t}=\left[f(\mathbf{x}^{t},t)-g^{2}(t)\nabla_{\mathbf{x}^{t}}\log p(\mathbf{x}^{t}|t)\right]\mathrm{d}t+g(t)\,\mathrm{d}W.$ (7)
This backward process is approximated using a deep neural network $\nabla_{\mathbf{x}^{t}}\log p_{\bm{\theta}}(\mathbf{x}^{t}|t)=s_{\bm{\theta}}(\mathbf{x}^{t},t)$. The learned backward SDE can be recast into a probability flow ODE (Maoutsa et al., 2020; Song et al., 2021), which in turn can be interpreted as a continuous normalizing flow (Chen et al., 2018) which facilitates efficient sampling and enables sample likelihood evaluation.
2.5 Implicit Transfer Operator Learning
Implicit Transfer Operator (ITO) Learning (Schreiner et al., 2023) is a framework for learning surrogate models of the transition density, $p(\mathbf{x}_{N\tau}|\mathbf{x}_{0})$, from MD data, where $N$ is an arbitrarily large integer. ITO leverages that the transfer operator framework allows us to express the transition density as
$p(\mathbf{x}_{t+N\tau}|\mathbf{x}_{t})=\sum_{i=1}^{\infty}\lambda_{i}^{N}(\tau)\psi_{i}(\mathbf{x}_{t})\phi_{i}(\mathbf{x}_{t+N\tau}),$ (8)
where $\psi_{i}$ and $\phi_{i}$ are independent of $\tau$ and $N$. See Appendix A.1 for a complete derivation.
This decomposition inspired a strategy to learn a conditional generative model $\mathbf{x}_{t+N\tau}\sim p_{\bm{\theta}}(\mathbf{x}_{t+N\tau}|\mathbf{x}_{t},N)$ by sampling tuples $(\mathbf{x}_{t_{i}},\mathbf{x}_{t_{i}+N_{i}\tau},N_{i})$ and training a generative model to mimic the empirical transition density at multiple time-horizons, $N\tau$. Here, we approximate ITO models with a conditional denoising diffusion probabilistic model (cDDPM) of the form
$p(\mathbf{x}^{0}_{t+N\tau}|\mathbf{x}_{t},N)\equiv\int p(\mathbf{x}^{0:T}_{t+N\tau}|\mathbf{x}_{t},N)\,\mathrm{d}\mathbf{x}^{1:T},$ (9)
where $\mathbf{x}_{1:T}$ are latent variables of the same dimension as our output, and follow a joint density describing the backward diffusion process, eq. 7, and $\mathbf{x}^{T}\sim\mathcal{N}(0,\mathbb{I})$. We denote diffusion time in Diffusion Models using superscripts, while physical time is represented using subscripts. The conditional sample likelihood is given by
$\ell(\mathbf{I};{\bm{\theta}})\equiv\prod_{i\in\mathbf{I}}p_{\bm{\theta}}(\mathbf{x}^{0}_{t_{i}+N_{i}\tau}\mid\mathbf{x}_{t_{i}},N_{i})$ (10)
where $\mathbf{I}$ is a list of generated indices $i$ specifying a time $t_{i}$ and a time-lag ($\tau$) integer multiple $N_{i}$, associating two time-points in a MD trajectory of length $M\tau$, $\mathbf{x}=\{\mathbf{x}_{0},\mathbf{x}_{\tau},\dots,\mathbf{x}_{(M-1)\tau}\}$.
Training is performed by optimizing an approximation of the variational bound of the log-likelihood (Ho et al., 2020),
$\mathcal{L}({\bm{\theta}})=\mathbb{E}_{i\sim\mathbf{I},\epsilon\sim\mathcal{N}(0,\mathbb{I}),t_{\mathrm{diff}}\sim\mathcal{U}(0,T)}\left[\lVert\epsilon-\hat{\epsilon}_{\bm{\theta}}(\widetilde{\mathbf{x}}^{t_{\mathrm{diff}}}_{t_{i}+N_{i}\tau},\mathbf{x}_{t_{i}},N_{i},t_{\mathrm{diff}})\rVert_{2}\right],$ (11)
where $\widetilde{\mathbf{x}}^{t_{\mathrm{diff}}}_{t}=\sqrt{\bar{\alpha}^{t_{\mathrm{diff}}}}\mathbf{x}_{t}+\sqrt{1-\bar{\alpha}^{t_{\mathrm{diff}}}}\epsilon$, with $\bar{\alpha}^{t_{\mathrm{diff}}}=\prod_{j}^{t_{\mathrm{diff}}}(1-\beta_{j})$ and $\beta_{j}$ is the variance of the forward diffusion process at diffusion time, $j$. $\hat{\epsilon}_{\bm{\theta}}(\cdot)$ is one of the two architectures presented by Schreiner et al. (2023). Following Arts et al. (2023) we express the score as
$s_{\bm{\theta}}(\mathbf{x}^{t_{\text{diff}}}_{t+N\tau},\mathbf{x}_{t},N_{i},t_{\text{diff}})=-\frac{\hat{\epsilon}_{\bm{\theta}}(\widetilde{\mathbf{x}}^{t_{\mathrm{diff}}}_{t_{i}+N_{i}\tau},\mathbf{x}_{t_{i}},N_{i},t_{\mathrm{diff}})}{\sqrt{1-\bar{\alpha}^{t_{\mathrm{diff}}}}}.$ (12)
3 Boltzmann Priors for Implicit Transfer Operator Learning
ITO training requires extensive, unbiased MD simulations to capture the statistical distribution of rare events. This data-intensive requirement can hinder ITO’s practical implementation. Furthermore, models trained on off-equilibrium simulations, which may exhibit non-representative statistics, can lead to inaccurate predictions, compromising their utility in downstream applications. In practice, however, both unbiased off-equilibrium simulations and information about the equilibrium distribution from potentially biased simulations (Hénin et al., 2022) can be cheaply generated, and both encode information about the dynamical behavior of molecular systems. We introduce Boltzmann Priors for Implicit Transfer Operator (BoPITO), as a learning paradigm for ITO models that leverage available information about the equilibrium distribution.
BoPITO uses pre-trained models of the equilibrium distribution to improve ITO learning in three ways. First, it helps ensure broad sampling of $\Omega$ proportional to the equilibrium distribution for subsequent MD simulations yielding information about the transition density $p(\mathbf{x}_{\tau}\mid\mathbf{x}_{0})$ from across $\Omega$. Second, we use it to fix the stationary part of the learned transition density, boosting the sample efficiency when learning models of molecular dynamics. Third, it imposes an inductive bias of long-time dynamical behavior allowing for recovery of unbiased Boltzmann distribution for long-time horizons. Fourth, using BoPITO, we introduce a novel tunable sampling protocol interpolating ITO models trained on off-equilibrium simulation data and an unbiased equilibrium distribution.
3.1 Efficient data generation
Training data for ITO models consists of several independent unbiased MD simulations. Following the adaptive sampling strategy (Bowman et al., 2010; Doerr & De Fabritiis, 2014; Viguera Diez et al., 2024; Betz & Dror, 2019), used extensively in the molecular dynamics simulation community, we use a pre-trained BG to generate initial conditions to simulations ensuring broad sampling across $\Omega$ proportional to the Boltzmann distribution. As long as these trajectories reach a ‘local equilibrium’ we can in principle recover an unbiased model of the molecular dynamics (Nüske et al., 2017), albeit without relying on running one or a few very long simulations to reach the global equilibrium.
3.2 Long-term dynamics inductive bias for ITO
One important application of ITO is to allow for one-step sampling of long-time-scale dynamics. However, real datasets often contain a limited number of effective samples for long time-scales, leading to potential biases in models. To mitigate this issue, we propose separating the equilibrium contribution from the time-dependent components:
$p(\mathbf{x}_{t+N\tau}|\mathbf{x}_{t})=\mu(\mathbf{x}_{t+N\tau})+\sum_{i=2}^{\infty}\lambda_{i}^{N}(\tau)\phi_{i}(\mathbf{x}_{t+N\tau})\psi_{i}(\mathbf{x}_{t}),$ (13)
where we have used that $\lambda_{1}=1$ and its corresponding eigenfunctions are $\phi_{1}=\mu$ and $\psi_{1}=\mathbf{1}$. Additionally, we introduce a decay in the time-dependent component, creating an inductive bias that asymptotically samples from an available equilibrium model for long-term dynamics. Using the spectral decomposition of the transition density (eq. 13), we choose the score model as
$s_{\bm{\theta}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}},{\bm{\theta}}),$ (14)
where $s_{\mathrm{eq}}(\mathbf{x}^{t}_{\text{diff}},t_{\text{diff}})$ is the score of a pre-trained surrogate of the equilibrium distribution model, $10<\hat{\lambda}<$ 1 is a hyper-parameter and $s_{\mathrm{dyn}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}},{\bm{\theta}})$ accounts for the time-dependent components. As $N\rightarrow\infty$, $s_{\mathrm{eq}}$ dominates and the model samples from the equilibrium model, see Appendix A.2 for an example. By interpreting the score field as a velocity field, we can reformulate the DM as a continuous normalizing flow which we can use for Metropolized sampling of the unbiased equilibrium distribution (Klein et al., 2023). The proposed factorization is principled, corresponding to a separation of the score of the transition density into a stationary and dynamic part, where the first part is considered known, see Appendix A.3 for details. In practice, we first train $s_{\mathrm{eq}}$ (if not provided) using equilibrium data. Then we train $s_{\mathrm{dyn}}$ with unbiased, possibly off-equilibrium, MD data while keeping $s_{\mathrm{eq}}$ fixed.
3.3 BoPITO interpolators
A significant issue in MD simulations is the unbiased sampling of transitions over high free-energy barriers, e.g. channels in $\mu(\mathbf{x})$ with very little probability mass, typically coinciding with mixing on $\Omega$. This difficulty leads to an inability to predict time-correlation statistics for large $N\tau$. With BoPITO we can define an interpolation between models trained on off-equilibrium simulations and the equilibrium distribution. In this manner, we can approximate dynamics not seen explicitly in the data. For a given maximum training lag, $N_{\text{max}}$, we define a BoPITO interpolator as a model with the following score function:
$s_{\bm{\theta}}(\mathbf{x}^{t_{\text{diff}}},\mathbf{x}_{t},N_{\text{int}},t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N_{\text{int}}}s_{\mathrm{dyn}}(\mathbf{x}^{t_{\text{diff}}},\mathbf{x}_{t},N_{\text{max}},t_{\text{diff}}),$ (15)
with $N_{\text{int}}\geq N_{\text{max}}$. This defines an interpolator because $N_{\text{int}}=N_{\text{max}}$ generates samples from the model distribution with maximum lag and $N_{\text{int}}\rightarrow\infty$ generates samples from the equilibrium model.
Inspired by methods in molecular biophysics (Kolloff & Olsson, 2023; Olsson et al., 2017; Bottaro & Lindorff-Larsen, 2018; Salvi et al., 2016), we propose to choose the interpolation parameter, $N_{\text{int}}$, as the most consistent with an unbiased dynamical observable, such as experimental data. That is, for a lag $N>N_{\text{max}}$ and an unbiased dynamic observable $O^{*}_{N}$, we choose
$N_{\text{int},N}=\operatorname*{arg\,min}_{N_{\text{int}}}|O^{*}_{N}-O_{N_{\text{int}}}|,$ (16)
where $O_{N_{\text{int}}}$ is the dynamic observable estimated with samples from the BoPITO interpolator. This way BoPITO can integrate off-equilibrium and equilibrium MD with experimental data. In practice, we find the interpolator to generate some high-energy states. However, we can alleviate the high-energy structures by alternating long-lag interpolation steps with short-lag non-interpolation steps for local relaxation.
4 Results
For detailed parameters of the experiments below, we refer to Appendix A.4.
4.1 Systems
Prinz potential is a 1D potential commonly used for benchmarking MD sampling methods (Prinz et al., 2011). We set the observable functions, $a$ and $b$, eq. 3, to be the identity function for computing dynamic observables. For details, see Appendix A.5.
Alanine Dipeptide is a small peptide with 22 atoms. We use publicly available data from Dibak et al. (2022), containing $1\,\mu\mathrm{s}$ simulation time split in $20$ trajectories. Simulation is performed in an implicit solvent with $2\,\mathrm{fs}$ integration time-step, and data is saved every $1\,\mathrm{ps}$. We choose,
$a(\mathbf{x})=b(\mathbf{x})=\begin{bmatrix}\sin{\phi}(\mathbf{x})\\ \cos{\phi}(\mathbf{x})\end{bmatrix},$ (17)
where $\phi(\mathbf{x})$ is a torsion angle involved in the slowest transition in the system. For details, see Appendix A.6.
4.2 Boltzmann priors for training data generation
Modeling the transition density over $\Omega$ requires observing transitions across the state space. In the usual setting, only one or a few initial conditions are available when data collection starts. Here, we explore the case where a BG is available before data collection and compare it against the baseline, where only one initial condition is known.
We use a BG trained on equilibrium data of the Prinz potential to sample the initial conditions of our training trajectories. We also generate trajectories using a single starting point (crystal, $\mathbf{x}=0.75$) and compare their performance with long simulations (Figure 2). We compare the two different data generation strategies using the metric, $|\Delta\text{correlation}|$, which measures the absolute difference of the time-correlation function (dynamic observable) compared to the MD ground truth (Appendix A.7). We find that the performance of models using Boltzmann priors for data generation is superior for different lag times (Figure 2a) and number of trajectories (Figure 2b). The gap between BG and the crystal baseline increases with lag-time, and as expected decreases with the number of generated trajectories.
Figure 2: Absolute difference in correlation (lower is better) of models trained on trajectories initialized on samples from a Boltzmann Generator (BG) and a single structure (crystal, $\mathbf{x}=0.75$) for the Prinz Potential under direct sampling. The former presents superior performance for different lag times (number of trajectories = $50$) (a) and number of trajectories (b). The shaded areas correspond to $95\ \%$ confidence interval.
4.3 BoPITO efficiently samples long-term dynamics in a low-data context
Fixing the equilibrium contribution to the score field effectively reduces the number of parameters that need to be estimated. To test whether this prior information manifests as an improved sample efficiency we compared ITO and BoPITO models against each other with varying sizes of training data.
We find that the BoPITO models achieve a higher accuracy for long-term dynamics compared to ITO models when data is scarce, as the equilibrium distribution is known a priori and does not need to be learned from simulation data (Figure 3). The inductive bias in BoPITO models enables them to learn long-term dynamics, even in scenarios where ITO models fail. For the Prinz Potential, we find that while ITO suffers from poor performance modeling long-term dynamics when data is scarce (Figure 3a), BoPITO models (Figure 3b) accurately capture long-term dynamics without worsening the performance on short and medium timescales. The results for Alanine Dipeptide mirror those of the Prinz potential: we find that an order of magnitude more data is needed to train an ITO (Figure 3c) to the same accuracy as a BoPITO model trained on the same data (Figure 3d). Moreover, we show in Appendix A.8 that the energies of configurations sampled from our prior Boltzmann Generator used in our experiments match closely the energies of MD samples.
Figure 3: Absolute difference in correlation (lower is better, one-step sampling) between ITO (left) and BoPITO (right) split into short, medium, and long time-scales against the number of training trajectories for the Prinz potential (top) and Alanine Dipeptide (bottom). The shaded areas correspond to a $95\ \%$ confidence interval.
4.4 Interpolating between models trained on off-equilibrium data and the Boltzmann distribution with experimental data
In practical settings, our MD data will be off-equilibrium, e.g. having sampled only one or a few of the relevant modes in the Boltzmann distribution in a given trajectory. Consequently, unless extensive data across the domain $\Omega$ can be collected, models based on such data will be biased. An alternative to collecting more simulation data is to use a multi-modal strategy where experimental data is used to fill the gaps left by simulation data and bridge to long time-scale dynamics (Salvi et al., 2016; Kolloff & Olsson, 2023). In this section we explore the potential of BoPITO interpolators to integrate dynamic observables with off-equilibrium simulation data to recover unbiased long-term dynamics.
Figure 4: Fitting an interpolator of a biased BoPITO model. Both the ITO and BoPITO models are trained on biased MD simulations of Alanine Dipeptide. We fit a BoPITO interpolator selecting the most consistent ensemble with the unbiased correlation function specified in Equation 16. Numbers with blue arrows specify the interpolation parameter, $N_{\text{int}}$.
We showcase BoPITO interpolators on Alanine Dipeptide. By removing the transitions between the modes of the Boltzmann distribution corresponding to the slowest process in the system, e.g. $\phi$ crossing 00 or $2$, we generate a biased simulation data, resembling a realistic scenario. We then train ITO and BoPITO models using these biased trajectories. For lags $>100$, we sample the BoPITO interpolator estimated by matching to an unbiased dynamic observable defined by eq. 22, allowing us to overcome a systematic error in the correlation function observed in the biased MD data and for ITO (Figure 4).
Beyond reproducing the provided dynamic observable, the interpolator also demonstrates a remarkable ability to capture the underlying microscopic dynamics (Figure 5). Even with excellent agreement, the generated ensembles only slightly overestimates density in the transition state region. However, this effect could easily be alleviated by annealing a short MD simulation to the interpolation as was recently shown (Viguera Diez et al., 2024).
5 Related work
Sampling the Boltzmann distribution
Apart from Boltzmann Generator-based approaches, there are a number of traditional ways to draw statistical samples from the Boltzmann distribution of molecular systems. Prominent examples include, molecular dynamics or Markov Chain Monte Carlo simulations coupled with enhanced sampling strategies (Hénin et al., 2022; Kamenik et al., 2022), including replica-based approaches (Earl & Deem, 2005; Sidler et al., 2016; Pasarkar et al., 2023), conformational flooding (Grubmüller, 1995), meta-dynamics (Laio & Parrinello, 2002), and umbrella sampling (Torrie & Valleau, 1977), in particular when paired up with machine-learned collective variables (Chen & Ferguson, 2018; Wang et al., 2019b; Herringer et al., 2023; M. Sultan & Pande, 2017). The success of these approaches relies on choosing the right mechanism to enhance sampling across high free energy barriers, or low probability regions, for any given case (Carter et al., 1989). Finding such mechanisms typically involves substantial manual engineering of collective variables and other hyper-parameters. Finally, there are transition path sampling approaches (Dellago et al., 1998; Bolhuis et al., 2002), which are particularly powerful when combined with reinforcement learning (Jung et al., 2023) or deep generative priors (Plainer et al., 2023).
Latent space simulators and Coarse graining
Classical (fine-grained) molecular dynamics simulation can be coarse-grained beyond the Born-Oppenheimer approximation, such that atomic nuclei are merged together into ‘beads.’(Noid, 2023) This strategy, in principle enables faster simulations due to the smaller number of particles and the acceleration of kinetics caused by the coarse-graining (CG) operation (Nüske et al., 2019). CG models can be estimated to closely approximate the thermodynamics of the corresponding fine-grained system through the optimization of two equivalent variational bounds (Noid et al., 2008; Lyubartsev & Laaksonen, 1995; Ercolessi & Adams, 1994). A relaxation of this bound was recently proposed to also allow for probabilistic reconstruction of fine-grained configurations (Chennakesavalu et al., 2023). These bounds have been used extensively to build CG force-field models (Husic et al., 2020; Wang et al., 2019a; Majewski et al., 2023; Charron et al., 2023) including implicit solvation models (Chen et al., 2021; Katzberger & Riniker, 2024), using deep neural networks due to their ability to capture multibody terms (Wang et al., 2021) to ultimately accelerate the prediction of equilibrium properties of molecular systems. Similarly, the development of ‘latent space simulators’ where a learned, typically low-dimensional, latent space equipped either with a propagator (Sidky et al., 2020; Chennakesavalu et al., 2023; Wang et al., 2024), or not (Wang & Gómez-Bombarelli, 2019), is learned to enable efficient simulation. These approaches, in general aim to accelerate molecular simulations akin to BoPITO, yet due to the CG operation, the molecular dynamics (kinetics) will be accelerated, and detailed knowledge of the unbiased dynamics are needed to correct this (Nüske et al., 2019; Crommelin & Vanden-Eijnden, 2011).
Figure 5: BoPITO can incorporate unbiased dynamic observables to correct a model trained on biased data. Rows of increasing time-lag (from top to bottom). Contour plots correspond to a BoPITO interpolator. The first column shows conditional transition densities projected onto the torsion angles $\phi$ and $\psi$ (inset). The black cross indicates the initial condition. The second and third columns show marginal distributions of $\phi$ and $\psi$, respectively.
Transfer Operator surrogates
Analysis of MD data often involves building transfer operator surrogates such as Markov state models (MSM) (Schütte et al., 1998; Prinz et al., 2011; Swope et al., 2004; Husic & Pande, 2018), time-lagged independent components analysis (Molgedey & Schuster, 1994; Ziehe & Müller, 1998; Pérez-Hernández et al., 2013), Markov field models or dynamic graphical models, (Olsson & Noé, 2019; Mardt et al., 2022; Hempel et al., 2022), VAMPnets (Mardt et al., 2018; Wu & Noé, 2019), or observable operator models (Wu et al., 2015). Markov state models are time-space discrete approximations of the transfer operator and Deep Generative MSM (Wu et al., 2019) and VAMPnets (Mardt et al., 2018) learn the space discretization through deep neural networks. Dynamic graphical models or Markov field models represent a time-space discrete approximation of the transfer operator injecting a (conditional) independence assumption of molecular subsystems, when modeling the transition probability, leading to better scaling for systems with poor time-scale separation (Olsson & Noé, 2019). Apart from ITO (Schreiner et al., 2023) several other deep generative approaches for modeling the transition density of molecular dynamics have recently proposed. Timewarp where a normalizing flow is used to encode the transition density (Klein et al., 2023) with limited transferability and metropolized sampling is used to access asymptotically unbiased equilibrium samples (Hastings, 1970). Score dynamics use a DDPM to model the displacements of an initial configuration towards a time-lagged one, achieving picosecond time-steps simulation and limited transferability (Hsu et al., 2024). However, unlike in the context of MSMs (Trendelkamp-Schroer et al., 2015; Rosta & Hummer, 2014; Wu et al., 2014; 2016), there are no deep generative transition density surrogates available leveraging available information about the equilibrium distribution — BoPITO is one such method.
6 Limitations and future work
Choice of hyper-parameter $\hat{\lambda}$
The hyper-parameter $\hat{\lambda}$, defines a global relaxation or mixing time-scale of the dynamics after which the model is guaranteed to sample equilibrium. In Appendix A.9 we describe a protocol to determine a bound for this parameter. However, developing a similar protocol for a transferable BoPITO model would likely require modifications to accommodate the diversity of global relaxation time-scales across different systems.
No Chapman-Kolmogorov Guarantee
When training on biased or off-equilibrium data where we rely on establishing ergodicity through interpolation we cannot guarantee self-consistency of the dynamics in the Chapman-Kolmogorov sense.
Surrogate model
BoPITO inherits the current limitations of ITO, such as generalization over chemical space and thermodynamic variables, and scaling. Furthermore, current models cannot guarantee unbiased sampling dynamics for non-equilibrium ensembles, which would require closed-form expressions for the target path probabilities.
7 Conclusion
We introduce Boltzmann Priors for Implicit Transfer Operator Learning (BoPITO), a framework to enhance ITO learning in two ways. First, a broad sampling of configuration space is used to initialize short off-equilibrium MD simulations. Second, we parameterize the transition density as an interpolation towards a pre-trained Boltzmann Generator, improving sample efficiency by an order of magnitude. BoPITO is a principled approach to embedding prior knowledge of the stationary distribution of Markovian dynamics as an inductive bias for long-term dynamical behavior. The approach enables interpolation between models trained on off-equilibrium data and the equilibrium distribution, and we can recover accurate models of unseen dynamics when informed by unbiased observables. Consequently, BoPITO is the first method to allow for the integration of multiple sources of information into the generation of deep generative surrogates of molecular dynamics.
Appendix A Appendix
Supporting information for Equilibrium priors and uncertainty estimation for deep generative models of molecular dynamics
Juan Viguera Diez1,2, Mathias Schreiner^1, Ola Engkvist1,2, and Simon Olsson1∗
A.1 Eigen decomposition of the Transfer Operator and transition density
The two-fold composition of the transfer operator, eq. 5, acting on an initial $\mu$-weighted density, $\rho$, is
$[T^{2}_{\Omega}(\tau)\circ\rho](\mathbf{x}_{t+2\tau})$ $=\sum_{i=1}^{\infty}\lambda_{i}(\tau)\left\langle\sum_{j=1}^{\infty}\lambda_{j}(\tau)\langle\rho|\phi_{j}\rangle\ \psi_{j}|\phi_{i}\right\rangle\ \psi_{i}(\mathbf{x}_{t+2\tau})$
$=\sum_{i=1}^{\infty}\lambda_{i}(\tau)\left(\sum_{j=1}^{\infty}\lambda_{j}(\tau)\langle\rho|\phi_{j}\rangle\ \langle\psi_{j}|\phi_{i}\rangle\right)\ \psi_{i}(\mathbf{x}_{t+2\tau})$
$=\sum_{i=1}^{\infty}\lambda^{2}_{i}(\tau)\langle\rho|\phi_{i}\rangle\ \psi_{i}(\mathbf{x}_{t+2\tau}),$
where we have used the orthonormality of the eigenfunctions, that is $\langle\psi_{j}|\phi_{i}\rangle=\delta_{ij}$. Similarly,
$[T^{N}_{\Omega}(\tau)\circ\rho](\mathbf{x}_{t+N\tau})=\sum_{i=1}^{\infty}\lambda^{N}_{i}(\tau)\langle\rho|\phi_{i}\rangle\ \psi_{i}(\mathbf{x}_{t+N\tau}).$ (18)
The transition density can be re-written in terms of the spectral decomposition of the transfer operator by choosing the initial density, $p$, as a Dirac delta function $\delta_{\mathbf{x}_{t}}(\mathbf{x})$, that is $\rho(\mathbf{x})=\frac{\delta_{\mathbf{x}_{t}}(\mathbf{x})}{\mu(\mathbf{x})}$. Then,
$p(\mathbf{x}_{t+N\tau}|\mathbf{x}_{t})$ $=\mu(\mathbf{x}_{t+N\tau})\left[T^{N}_{\Omega}(\tau)\circ\frac{\delta_{\mathbf{x}_{t}}}{\mu}\right](\mathbf{x}_{t+N\tau})$
$=\mu(\mathbf{x}_{t+N\tau})\sum_{i=1}^{\infty}\lambda^{N}_{i}(\tau)\left\langle\ \frac{\delta_{\mathbf{x}_{t}}}{\mu}\middle|\phi_{i}\right\rangle\ \psi_{i}(\mathbf{x}_{t+N\tau})$
$=\sum_{i=1}^{\infty}\lambda^{N}_{i}(\tau)\left\langle\ \delta_{\mathbf{x}_{t}}(\mathbf{x})|\psi_{i}\right\rangle\ \psi_{i}(\mathbf{x}_{t+N\tau})\ \mu(\mathbf{x}_{t+N\tau})$
$=\sum_{i=1}^{\infty}\lambda_{i}^{N}(\tau)\psi_{i}(\mathbf{x}_{t})\phi_{i}(\mathbf{x}_{t+N\tau}).$
A.2 Decay of time-dependent score term
In Figure 6, we show the average time-dependent component of the score,
$s_{N}=\mathbb{E}_{i\sim\mathbf{I},t_{\mathrm{diff}}\sim\mathcal{U}(0,T)}\left[\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{t_{i}+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t_{i}},N,t_{\text{diff}},{\bm{\theta}})\right],$ (19)
of a trained model BoPITO model of the Prinz Potential with $\hat{\lambda}=0.994$. The model remains flexible for small lags but eventually decreases to 00, sampling from the equilibrium model.
Figure 6: Average time-dependent component of the score, eq. 19, of a trained model BoPITO model of the Prinz Potential.
A.3 The score of the transition probability and BoPITO
The score of the transition probability can be written as
$\begin{split}\nabla_{\mathbf{x}_{N\tau}}\log p(\mathbf{x}_{N\tau}|\mathbf{x}_{0},N)&=\sum_{i}^{\infty}\lambda_{i}(\tau)^{N}\frac{\psi_{i}(\mathbf{x}_{0})}{p(\mathbf{x}_{N\tau}|\mathbf{x}_{0},N)}\nabla_{\mathbf{x}_{N\tau}}\phi_{i}(\mathbf{x}_{N\tau})\\ &=\frac{\nabla_{\mathbf{x}_{N\tau}}\mu(\mathbf{x}_{N\tau})}{p(\mathbf{x}_{N\tau}|\mathbf{x}_{0},N)}+\sum_{i=2}^{\infty}\lambda_{i}(\tau)^{N}\frac{\psi_{i}(\mathbf{x}_{0})}{p(\mathbf{x}_{N\tau}|\mathbf{x}_{0},N)}\nabla_{\mathbf{x}_{N\tau}}\phi_{i}(\mathbf{x}_{N\tau})\\ &=\frac{\nabla_{\mathbf{x}_{N\tau}}\mu(\mathbf{x}_{N\tau})}{\mu(\mathbf{x}_{N\tau})}\frac{\mu(\mathbf{x}_{N\tau})}{p(\mathbf{x}_{N\tau}|\mathbf{x}_{0},N)}\\ &+\hat{\lambda}^{N}\sum_{i=2}^{\infty}\left(\frac{\lambda_{i}(\tau)}{\hat{\lambda}}\right)^{N}\frac{\psi_{i}(\mathbf{x}_{0})}{p(\mathbf{x}_{N\tau}|\mathbf{x}_{0},N)}\nabla_{\mathbf{x}_{N\tau}}\phi_{i}(\mathbf{x}_{N\tau}).\end{split}$
$\frac{\nabla_{\mathbf{x}_{N\tau}}\mu(\mathbf{x}_{N\tau})}{\mu(\mathbf{x}_{N\tau})}=\nabla_{\mathbf{x}_{N\tau}}\log\mu(\mathbf{x}_{N\tau})$ is the score of the Boltzmann distribution and can be modeled with the score of a Boltzmann Generator, $s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})$. $\frac{\mu(\mathbf{x}_{N\tau})}{p(\mathbf{x}_{N\tau}|\mathbf{x}_{0},N)}\rightarrow 1$ as $N\rightarrow\infty$ and can be modeled with a scalar neural network $1+\hat{\lambda}^{N}f_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})$. The last term can be modeled with $g_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})$. The corresponding score model is
$s_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})(1+\hat{\lambda}^{N}f_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})+\hat{\lambda}^{N}g_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})$
$=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}\left(s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})f_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})+g_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})\right).$
We can aggregate $f$ and $g$ to a single neural network component, $s_{\mathrm{dyn}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})$, to get
$s_{\bm{\theta}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{N\tau}^{t_{\text{diff}}},\mathbf{x}_{0},N,t_{\text{diff}})$ (20)
A.4 Experimental parameters
A.4.1 Boltzmann Priors for dataset generation
For different number of trajectories, $n$, we train $5000/n$ ITO models on $n$ Prinz potential trajectories of length 150. Trajectories do not overlap among different trainings. The maximum model lag is 100 steps.
A.4.2 BoPITO samples efficiently long-term dynamics in data-sparse scenarios
Prinz potential
[…]
Alanine Dipeptide
[…]
A.4.3 Interpolating between models trained on off equilibrium data and the Boltzmann distribution
We remove the transitions resembling the slowest process in the system, $\phi$ crossing 00 or $2$, to generate biased simulation data. When a transition occurs, we remove one frame before and after the transition and split the trajectory. We train both biased ITO and BoPITO models with a maximum model lag of $100$ on the resulting biased dataset. For lags $>100$, we sample the BoPITO interpolator fitted on the unbiased dynamic observable defined by eq. 22, see Figure 4. We perform interpolation by sampling with the score model in eq. 15 followed by one round of non-interpolation sampling with lag $100$ steps for local relaxation.
A.4.4 Architectural and training details
We use MB-ITO as the architecture of models for the Prinz potential, and SE3-ITO, an $SE(3)-$equivariant neural network, for Alanine Dipeptide, both introduced in Schreiner et al. (2023). We report architectural and training hyper-parameters in Table 1. Models are trained until convergence in the log-log loss plot.
Diffusion steps $500$
Noise schedule Sigmoidal
Batch size $2,097,152$
Learning rate $0.001$
Layers $3$
Embedding dimension $256$
Net dimension $256$
Optimizer Adam
Diffusion steps $1,000$
Noise schedule Polynomial
Batch size $1,024$
Learning rate $0.001$
Score layers $5$
Embedding layers $2$
n_features $64$
Optimizer Adam
Table 1: Architectural and training parameters of MB-ITO models (a) and SE3-ITO (b).
A.5 Prinz potential
The Prinz potential is a 1D potential commonly used for benchmarking MD sampling methods. The potential is defined as
$U(x)=4\left(x^{8}+0.8e^{-80x^{2}}+0.2e^{-80(x-0.5)^{2}}+0.5e^{-40(x+0.5)^{2}}\right).$ (21)
We generate trajectories using an Euler-Maruyama integrator using the library Deeptime (Hoffmann et al., 2021). We set the integrator time-step to $1\cdot 10^{-5}$, and the temperature, mass, and damping factor to 1. In Figure 7 we show histograms of the position of a particle after $N$ steps, starting from $0.75$, and in Figure 8 (a) we report an implied time-scales plot. We use the identity function to define a dynamic observable/correlation function of this system and show it in Figure 8 (b).
Figure 7: Conditional density of a simulation of the Prinz potential starting at $x=0.75$ for different lags (steps). Long-term dynamics approaches the Boltzmann distribution.
Figure 8: Implied time-scales (a) and dynamic observable under the identity function (b) for the Prinz Potential. Errors in (b) are smaller than the dots.
A.6 Alanine Dipeptide
Alanine Dipeptide is a small peptide with 22 atoms. We use publicly available data from Dibak et al. (2022), containing a total of $1\ \mu s$ simulation time split in $20$ trajectories. Simulation is performed in an implicit solvent with $2\ fs$ integration time-step and data is saved every $1\ ps$. We choose,
$a(x)=b(x)=\begin{bmatrix}\sin{\phi}\\ \cos{\phi}\end{bmatrix},$ (22)
to define a dynamic observable of this system. The torsion angle $\phi$ is involved in the slowest transition observed in the simulation, see Figure 10. We combine these vectors taking the inner product for computing dynamic observables. We show this correlation function in Figure 9 (a), and implied time-scales plot in Figure 9 (b) and a histogram of the torsions $\phi$ and $\psi$ aggregating all simulation data in Figure 10.
Figure 9: Implied time-scales for Alanine Dipeptide computed with a Markov State Model on torsional angles $\phi$ and $\psi$ (a) and dynamic observable in eq. 22 (b). Errors in (b) are smaller than the dots.
Figure 10: Histogram of the torsional angles $\phi$ and $\psi$ of Alanine Dipeptide (insert). Data is aggregated among all trajectories in the dataset introduced in Dibak et al. (2022).
A.7 Metrics
We evaluate models computing differences in the dynamic observables introduced in section 4.1 w.r.t. long MD simulations,
$|\Delta\text{correlation}|_{N}=|O^{*}_{N}-O^{\text{model}}_{N}|,$ (23)
where $O^{*}_{N}$ is the normalized observable predicted by MD for lag $N$ and $O^{\text{model}}_{N}$ is the model’s prediction. See Appendix A.10 for details on normalization. We report the average difference over different training runs.
A.8 Energies for Alanine Dipeptide
In Figure 11 we observe a remarkable agreement in energy densities of samples generated with our Boltzmann Generator and MD. GBSAOBForce is the implicit solvent model energy.
Figure 11: Energy and energy components of samples generated with a Boltzmann Generator (BG) and MD.
A.9 Fitting $\hat{\lambda}$
The hyperparameter $\hat{\lambda}$ controls the time-scale at which the BoPITO model transitions to sampling the equilibrium model. If $\hat{\lambda}$ is too small, the model may prematurely relax to equilibrium, limiting its ability to capture non-equilibrium dynamics. Conversely, if $\hat{\lambda}$ is too large, the benefits of the BoPITO framework could diminish. Therefore, careful tuning of $\hat{\lambda}$ can be crucial for optimal performance.
As illustrated in Figure A.9, grid-search hyperparameter tuning can select an appropriate $\hat{\lambda}$ for a single-system model. Too small values of $\hat{\lambda}$ can hinder the model’s ability to capture long-time-scale dynamics, leading to increased loss. We recommend choosing the smallest $\hat{\lambda}$ that yields a plateau in the $\hat{\lambda}$-loss curve (elbow rule), as demonstrated in Figure A.9 (b). However, practitioners should be cautious about sampling longer lags than the implied time-scale defined by $\hat{\lambda}$ if they cannot guarantee the system relaxes to equilibrium for those time-scales.
Figure 12: Average validation loss against training epochs for the Prinz potential for different values of $\hat{\lambda}$ (a) and average final loss versus $\hat{\lambda}$ (b). Averages are taken w.r.t. 10 runs. Final losses are computed as the average between epochs $190$ and $200$. The greatest eigenvalue of the system under our simulation parameters is $0.994$.
A.10 Correlation function normalization
We subtract the mean and divide by correlation at lag 00 to normalize our dynamic observables. Subtracting the mean guarantees that the correlation function asymptotically decays to 00, and we divide by the correlation at lag 00 so that correlation can be read as a fraction of correlation at time 00. The resulting normalized correlation function is
$\frac{E[\left(f(\mathbf{x}_{t})-E[f(\mathbf{x}_{t})]\right)\left(g(\mathbf{x}_{t+\Delta t})-E[g(\mathbf{x}_{t+\Delta t})]\right)]}{E[\left(f(\mathbf{x}_{t})-E[f(\mathbf{x}_{t})]\right)\left(g(\mathbf{x}_{t})-E[g(\mathbf{x}_{t})]\right)]}.$ (24)
When comparing different methods, we compute $E[f(\mathbf{x}_{t})]$, $E[g(\mathbf{x}_{t+\Delta t})]$ and $E[\left(f(\mathbf{x}_{t})-E[f(\mathbf{x}_{t})]\right)\left(g(\mathbf{x}_{t})-E[g(\mathbf{x}_{t})]\right)]$ using long MD simulations and use these same normalizing factors for all methods.